An isosceles triangle has two sides of equal length. The length of one of the eqaul sides is 2 more than three times the length of the third side. If the perimeter is 46 m, find the length of each side.
I need to translate this sentence into an equation and then solve

Answer by rfer(16322)

2(3x+2)+x=46
6x+4+x=46
7x=42
x=6 one side
3x+2=20 two sides